find the slope of a line perpendicular to the the line that passes through the following points.
(3,9) and (7,15)

Answer by jim_thompson5910(35256)

First find the slope of the line through the two points.


m = (y2-y1)/(x2-x1)


m = (15-9)/(7-3)


m = 6/4


m = 3/2


So the slope of the line that passes through the two points is 3/2


The perpendicular slope is simply the negative reciprocal of that slope. Basically, flip the fraction and the sign to get the perpendicular slope.


So the perpendicular slope is -2/3